

********************************************************************
UARS WEEKLY STATUS REPORT - 31 MAY 2002
********************************************************************
***********
SOLSTICE
***********
SOLSTICE resumed full time operations this week, performing our complete
suite of solar observations.
***********
SUSIM
***********
SUSIM resumed operatons on 28 May. Since then, daily mid-resolution
and MgII solar scans have been successfully performed during periods
of predicted UARS telemetry.
***********
ACRIM
***********
The ACRIM2 Science Team is currently studying a command scenario designed
to stimulate the bootstrap circuit in its power supply module. We believe a
procedure having the same effect on ACRIM2's power supply as the TMON event
of August 2001 might be successful.
***********
PEM
***********
PEM wishes to thank the SUSIM team for producing UARS attitude files.
The PEM team was able to up-load these data to CDHFB:: and generate the PEM
O/A subset information for its processing of science data.
PEM support of the UARS Traceability Mission for the TIMED mission has
been organized with the GUVI team. PEM data was collected from April 18
through April 24. During this time period, there were two spacecraft contacts.
Early PEM data from the contact times showed no aurora. No significant aurora
appeared at the predicted highest invariant latitudes; however, lower invariant
latitude passes did show significant aurora. It appears as though magnetic
activity increased at times when UARS was not at its peak invariant latitudes
and some of the particle precipitation activity was captured by the PEM
instrumentation. PEM wishes to thank Bob Neff and Rich Beck at GSFC for their
efforts in retriving as much of the PEM data as possible.
The PEM specific orbit/attitude and high-energy particle data processing
has been completed. A selected set of PEM images showing precipitating
electrons and protons has been produced and is available at
"http://pemrac.space.swri.edu/guvi-pem.html".
PEM has provided input to the "Operations Agreement (OA) Between the
Packet Processor-Automation (Pacor-A) and Particle Environment Monitor (PEM)
for Upper Atmosphere Research Satellite (UARS)". This document is part of the
Consolidated Space Operations Contract.
PEM is working with the GUVI team to identify a second time period for
coordinated measurements. Early predictions show the best time of conjunction
is just after the UARS YAW at the end of June; however, the spacecraft power
my not be sufficient to operate PEM. The orbital conditions are different
from the first period and the TIMED-UARS orbits are nearly perpendicular at
this time. More investigation into what electron spectra are predicted needs
to be performed in order to determine if it is feasible to operate a GUVI-PEM
campaign after the June YAW.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nolthing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations to support re-engineering of the UARS
Mission Operations Concept to develop capabilities for the
Science Traceability Mission.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 18 May 2002 (Orbit 58415, GMT Day 138) through
24 May 2002 (Orbit 58520, GMT Day 144).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 140/2323 58460 POWER OFF
SUSIM 142/0020 58475 POWER OFF
PEM 142/0020 58475 POWER OFF
HRDI 142/1400 58484 LOW POWER MODE
SOLSTICE 144/0251 58507 TO DAY-ONLY MODE
A Forward-to-Reverse Yaw Around maneuver (number 109) was successfully
executed in Orbit 58500 beginning 143/15:34:35Z and completing at
143/16:11:45Z on 23 May 2002. The SOLSTICE instrument was in day only
and all other instruments were in their off or standby mode (including
HRDI and PEM). Due to the SA position at 269 degrees, the effective
night length of the spacecraft after the maneuver was approximately
57.6 minutes. Post maneuver analysis of data indicate that during the
Yaw Around, the end-of-night load bus voltage was 24.8 volts and the
average depths of discharge (DODs) on Batteries 2 and 3 were
approximately 20.50 and 19.89 percent respectively. The next (reverse-
to-forward) YAW is scheduled for 21 June 2002 (DOY 172).
The beta angle ranged between the angles of 21.9 degrees to 02.0
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 34 degrees on 08 June 2002 (DOY 159).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.8 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.96 to 0.98 degrees C. Additional
battery performance data for each battery on 17 May and 24 May 2002
are provided in Appendix D.
The clock error ranged from -10.2 to 11.0 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 06
HALOE = 01
SUSIM = 01
PEM = 01
HRDI = 01
SOLSTICE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 3
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 01/333 - 2287.496855
Transponder B: 02/043 - 2287.498879
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT is in the process of implementing some minor changes to the
Eclipse system to help optimize operations. All of the existing
documentation is also in the process of being updated to reflect the
new ground system nomenclature. A few new SOPs have yet to be
written, while the handful that have already been written need to be
finalized and signed off on. A new release of the Eclipse software is
scheduled for the end of the month, which should take care of most of
the changes needed.
MEETINGS
The FOT attended a UARS/ERBS Meeting on 22 May 2002 to discuss the
integration of the ERBS satellite into the UARS control center.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 21 June 2002 (DOY 172). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 24 May 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
58460 140/2323 02-181 HALOE POWER OFF
58468 141/1220 02-182 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
58475 142/0020 02-183 SUSIM POWER OFF
58475 142/0020 02-184 PEM POWER OFF
58484 142/1400 02-185 HRDI LOW POWER MODE
58488 142/1952 02-186 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
58499 143/1359 02-187 OBC VT LEVEL #6
58499 143/1439 02-189 MACS TAM TO MONITOR MODE
58500 143/1606 02-188 OBC YAW TO REVERSE FLIGHT
58502 143/1822 02-187 OBC VT LEVEL #5
58507 144/0251 02-191 SOLSTICE TO DAY-ONLY MODE
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 24 May 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
___________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 24 May 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
58468 141/1225 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
58488 142/1957 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 24 May 2002
17 May 2002 - GMT Day 137 Beta = 23.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.58/-4.52 +00.0/+00.0 0.0 0.000 0.0
2 +11.2/ -51.6 +5.15/+3.87 +27.7/-13.5 25.1 1.030 20.6
3 +22.4/ -33.6 +3.87/+2.92 +26.7/-13.5 25.1 1.020 20.1
24 May 2002 - GMT Day 144 Beta = 02.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.68/-4.62 +00.0/+00.0 0.0 0.000 0.0
2 +11.2/ -72.8 +5.47/+4.51 +27.2/-13.6 23.8 1.020 20.7
3 +22.4/ -33.6 +4.19/+3.23 +26.8/-13.6 23.8 1.010 20.4
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
58499 143/1359 L5 -> L6
58502 143/1822 L6 -> L5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 24 May 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-026 FOT 58446 140/0105 14170 Generic Late Acquisition
02-029 FOT 58377 135/1050 24334 Late Acquisition
02-030 FOT 58464 141/0533 24348 Equipment failover
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-026 FOT See List May 01 N/A Generic Late Acquisition
02-004 SOLSTICE 56883 036/0050 N/A Solstice in Wrong Mode
02-023 FOT 57767 094/2027 24285 Early LOS on OMNI Pass
02-024 FOT 57783 095/2209 24285 SOLSTICE INOPS in Day
02-027 FOT 58202 123/1931 24324 Late Acquisition
02-028 FOT 58277 128/2004 24331 Late Acquisition - 4:02
02-029 FOT 58377 135/1050 24334 Late Acquisition
02-030 FOT 58464 141/0533 24348 Equipment failover
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-029 FOT 58377 135/1050 24334 Late Acquisition
02-030 FOT 58464 141/0533 24348 Equipment failover